PI: Rioja/Sanghera
Title: CAL - 0016+731

Session: Oct 1996 (session 3)
Frequency: 6 cm
Mode: A (dual polarisation)
Obs time (UT): 295 EU 17:30(21/10)-18:30(21/10)
Stations: EbJbNtOnShSmToUrWb
Processor: Bonn
EU Mk3 Tapes/St.: 1


To give feedback on this experiment, click here

Experiment Feedback

Onsala (J. Conway, A. Polatidis at Wed Oct 23 04:52:32 EDT 1996) Minor failures
First scan lost due to error in the "label" command. The remaining scans succesful. Rain.

Effelsberg (D.Graham at Wed Oct 23 09:17:56 EDT 1996) Success

Effelsberg (D.Graham at Wed Oct 23 09:22:57 EDT 1996) Success
No known problems

Westerbork (tony foley at Thu Oct 24 03:48:50 EDT 1996) Success
Looks OK. Note that this was observed in singledish mode using telescope 6 and the prototype MFFE.

Torun (Roman Feiler at Thu Oct 24 17:07:07 EDT 1996) Minor failures
No known problems

Shanghai (Huang Xinyong at Fri Oct 25 03:16:39 EDT 1996) Minor failures
Clock offset = 0.3 micro sec. All obs took place, except data is invalid from 17:38:29 to 17:43:04 because tape vaccum failed. Weather was overcast.

Noto (Corrado Trigilio at Mon Oct 28 04:36:14 EST 1996) Success
No known problems. Good weather.

Jodrell (Alastair Gunn at Thu Nov 7 06:40:21 EST 1996) Success
No known problems. Tape IRAM0080 shipped to Bonn on 28/10/96 via Woodcroft.

Urumqi (Liu Xiang at Mon Nov 18 08:50:54 EST 1996) Failure
Was not been observed because of MK3 failed accidently.

Simeiz (Alexandre Volvatch at Tue Nov 19 04:00:37 EST 1996) Success
Clock offset 0.2 microsecond. No problems. Temp: 15 C. Press: 764 mm hg. Humid: 62 % and it was raining.